A ray of light in corn oil is incident striking with an angle of incidence of 40 degrees on water. Calculate the angle of refraction to the nearest degree


1
Expert's answer
2021-07-13T13:30:20-0400

"Gives"


"n_i=1.48,n_w=1.33\n,i=40\u00b0,r=?"

"n_isin\theta_i=n_wsin\theta_r"

"1.48\times sin40°=1.33 \times sin\theta_r"

"1.48\times0.642=1.33\times sin\theta_r"

"sin\theta_r=0.7145"

"\theta_r=46°"